OpenStudy (anonymous):

The minute hand on a watch is 5 mm long and the hour hand is 4 mm long. How fast is the distance between the tips of the hands changing at one o'clock? (Round your answer to one decimal place.)

OpenStudy (anonymous):

 Does this help? you can find a relationship between the angle and x through the law of cosines. and you also know rate at which both hands are moving so you can find the rate at which the angle is changing.

OpenStudy (anonymous):

but i cannot calculate the rate of angle.

OpenStudy (anonymous):

sure you can, the hour hand move at a rate of: \[\frac{360^o}{24 hours} =\frac{15^o}{hour} \] while the minute hand moves at a rate of: \[\frac{360^0}{hour}\] if the minute hand is moving towards the hour hand it must be?

OpenStudy (anonymous):

my bad, the hour hand is twice that...cause it makes one complete revolution in 12 hours. so it would be: \[\frac{30^o}{hour}\]